What is X+1/3+X+2/7=2
Answer by fractalier(6550)

Okay, from
x + 1/3 + x + 2/7 = 2
let's collect like terms and get
2x + (1/3 + 2/7) = 2
From here we can either combine fractions (probably easiest) or multiply through by a common denominator to clear all fractions...let's combine them...
1/3 + 2/7 = 7/21 + 6/21 = 13/21
so we have
2x + 13/21 = 2
now subtract 13/21 from both sides
2x = 2 - 13/21 = 42/21 - 13/21 = 29/21
now divide both sides by two and get
2x = 29/21
x = (1/2)(29/21)
x = 29/42
ugly answer, but the right one...